Consider the following data for the airline industry for July 2018 ​(EV=enterprise ​value, Book=equity book​ value). Suppose Alaska Air​ (ALK) has 123123 million shares outstanding. Estimate Alaska​ Air's share value using each of the five valuation​ multiples, based on the median valuation multiple of the other seven airlines shown.

Market Capitalization Enterprise Value (EV) EV/Sales EV/EBITDA EV/EBIT P/E Forward P/E Total Shares
Alaska Air (ALK) 7,286 8,207 1.02 5.1 6.7 7.7 12.5 123
Peer Data
American Airlines (AAL) 17,879 37,327 0.87 5.7 8.2 10.5 8.2
Delta Air Lines (DAL) 35,580 44,164 1.02 5.7 8 10.8 8.8
Hawaiian (HA) 2,023 2,057 0.73 3.4 4.1 5.8 7.2
JetBlue Airways (JBLU) 5,450 5,814 0.8 4.1 5.9 4.9 10.8
SkyWest (SKYW) 2,909 5,040 1.56 7.2 12.6 6.6 11.8
Southwest Airlines (LUV) 30,318 30,691 1.44 6.9 8.9 8.6 12.3
United Continental (UAL) 21,961 31,393 0.8 5.7 9.2 11.4 9.3
Source: CapitalIQ
Answer #1

Calculation of financial data of Alaska for calculating multiples

EV = Market Capitalization + (Debt - Cash)

EV of Alaska = 8207

Market Capitalization = 7286

Debt - Cash = EV - Market Capitalization = 8207 - 7286 = 921

EV/Sales = 1.02

Sales = EV / 1.02 = 8207 / 1.02 = 8046.08

EV/EBITDA = 5.1

EBITDA = EV / 5.1 = 8207 / 5.1 = 1609.22

EV/EBIT = 6.7

EBIT = EV / 6.7 = 8207 / 6.7 = 1224.93

Price per share = Market Capitalization / Total Shares

Price per share = 7286 / 123 = 59.24

Price per Share / Earning per share (P/E) = 7.7

Earning per share = Price per share / 7.7 = 59.24 / 7.7 = 7.69

Forward P/E = Price per share / Forward earning = 12.5

Forward Earning per share = Price per share / 12.5 = 59.24/12.5 = 4.74

Calculation of Median Multiples

Multiples EV/Sales EV/EBITDA EV/EBIT P/E Forward P/E
Peer Data
American Airlines (AAL) 0.87 5.7 8.2 10.5 8.2
Delta Air Lines (DAL) 1.02 5.7 8 10.8 8.8
Hawaiian (HA) 0.73 3.4 4.1 5.8 7.2
JetBlue Airways (JBLU) 0.8 4.1 5.9 4.9 10.8
SkyWest (SKYW) 1.56 7.2 12.6 6.6 11.8
Southwest Airlines (LUV) 1.44 6.9 8.9 8.6 12.3
United Continental (UAL) 0.8 5.7 9.2 11.4 9.3
Median 0.87 5.7 8.2 8.6 9.3

Calculation of Share Price

EV/Sales = .87

Sales = 8046.08

EV = .87 * Sales = .87 * 8046.08 = 7000.09

Debt - Cash = 921

Market Capitalization = EV - (Debt-cash) = 7000.09 - 921 = 6079.09

Price per share = Market Capitalization / Number of shares = 6079.09 / 123 = 49.42

EV/EBITDA = 5.7

EBITDA = 1609.22

EV = 5.7 * EBITDA = 5.7 * 1609.22 = 9172.55

Debt - Cash = 921

Market Capitalization = EV - (Debt-cash) = 9172.55 - 921 = 8251.55

Price per share = Market Capitalization / Number of shares = 8251.55 / 123 = 67.09

EV/EBIT = 8.2

EBIT = 1224.93

EV = 8.2 * EBIT = 8.2 * 1224.93 = 10044.43

Debt - Cash = 921

Market Capitalization = EV - (Debt-cash) = 10044.43 - 921 = 9123.43

Price per share = Market Capitalization / Number of shares = 9123.43 / 123 = 74.17

P/E = 8.6

Earning per share = 7.69

Price per share = 8.6 * Earning per share = 8.6 * 7.69 = 66.13

Forward P/E = 9.3

Forward Earning per share = 4.74

Price per share = 9.3 * Forward Earning per share = 9.3 * 4.74 = 44.08

Summary

Share Value using each of 5 valuation multiples

Shares Value based on EV/Sales Multiple = 49.42

Shares Value based on EV/EBITDA Multiple = 67.09

Shares Value based on EV/EBIT Multiple = 74.17

Shares Value based on P/E Multiple = 66.13

Shares Value based on Forward P/E Multiple = 44.08

Note : Excel formula (=MEDIAN) has been used to calculate median multiples